# Parametric Expensive Multi-Objective Optimization via Generative Solution Modeling

### v2

@@ Abstract
Many real-world applications require solving families of expensive multi-objective optimization problems (EMOPs) under varying operational conditions. This gives rise to parametric expensive multi-objective optimization problems (P-EMOPs) where each task parameter defines a distinct optimization instance. Current multi-objective Bayesian optimization methods have been widely used for finding finite sets of Pareto optimal solutions for individual tasks. However, P-EMOPs present a fundamental challenge: the continuous task parameter space can contain infinite distinct problems, each requiring separate expensive evaluations. This demands learning an inverse model that can directly predict optimized solutions for any task-preference query without expensive re-evaluation. This paper introduces the first parametric multi-objective Bayesian optimizer that learns this inverse model by alternating between (1) acquisition-driven search leveraging inter-task synergies and (2) generative solution sampling via conditional generative models. This approach enables efficient optimization across related tasks and finally achieves direct solution prediction for unseen parameterized EMOPs without additional expensive evaluations. We theoretically justify the faster convergence by leveraging inter-task synergies through task-aware Gaussian processes. Meanwhile, empirical studies in synthetic and real-world benchmarks further verify the effectiveness of our alternating framework.
@@ 1 Introduction
Expensive multi-objective optimization problems (EMOPs) are crucial in black-box optimization, aiming to identify optimal trade-off solutions among multiple conflicting criteria under limited evaluation budgets [parego-1, pareto-n]. Due to their focus on conflicting objective functions and costly evaluations, EMOP formulations have been applied to a plethora of real-world applications encompassing materials discovery [Merchant2023, beichen2024], drug discovery [shin2025offline, moo-gflownet], and robotics [ICRA-1, ICRA-2].
@@
To efficiently identify trade-off solutions, surrogate modeling has become essential for reducing the cost of evaluating expensive objectives. Among these, multi-objective Bayesian optimization (MOBO) stands out as the leading strategy. MOBO methods have been widely used for identifying finite sets of Pareto optimal solutions by scalarizing multiple objectives into single objectives [parego-1, sms-ego, parego-2] or by optimizing metrics such as expected hypervolume improvement [ehvi-1, ehvi-3, ehvi-4] and uncertainty reduction [uncertainty-1, uncertainty-2, uncertainty-3, uncertainty-4]. To move beyond these finite solution‑set approximations and directly model the entire trade‑off surface, recent efforts have extended MOBO to continuous Pareto‑manifold learning [moo-psl, moo-DA, moo-DM].
@@
However, in many real-world scenarios, optimization problems arise not as isolated instances but as families of related tasks under varying operational conditions. For example, turbine components must be optimized across different wind speeds [moo-gamut], and robotic control policies must be optimized for diverse terrains [wang2023softzoo] or damage conditions [nature]. These cases create fundamental challenges: the continuous range of operational conditions forms infinitely many related optimization problems, each demanding separate, costly evaluations under current MOBO methods. As operational conditions frequently shift [moo-context-ICRA], repeatedly performing full optimizations to evaluate trade-offs across multiple scenarios becomes prohibitively expensive and computationally infeasible. Therefore, handling families of related EMOPs requires new approaches that avoid excessive evaluation costs. To our best knowledge, little existing work has systematically addressed this parametric setting.
@@
This paper proposes the first parametric multi-objective optimizer to address this challenge by learning an inverse model. This inverse model can directly predict optimized solutions for any given task parameter, thereby bypassing the repeated costly evaluations mentioned above. Our approach alternates between two complementary phases: (1) acquisition-driven search guided by task-aware Gaussian processes, which explicitly leverages inter-task correlations to efficiently explore the solution space, and (2) generative solution sampling via conditional generative models, which learns high-quality solution distributions conditioned on task parameters and preferences. This dual strategy not only significantly enhances optimization efficiency across multiple related tasks but also finally builds an inverse model for predicting optimized solutions to previously unseen EMOPs without additional expensive evaluations.
@@
  • We propose the first inverse model to predict optimized solutions to each task query within the continuous task parameter space, thereby addressing the challenge of solving families of EMOPs.
@@ 2.1 Expensive Multi-Objective Optimization
Solution $\mathbf{x}^{(a)}\in\mathcal{X}$ is considered Pareto dominate another solution $\mathbf{x}^{(b)}\in\mathcal{X}$ if $\forall i\in[M]$, $f_{i}(\mathbf{x}^{(a)})\leq f_{i}(\mathbf{x}^{(b)})$ and $\exists\,i^{\prime}\in[M]$ such that $f_{i^{\prime}}(\mathbf{x}^{(a)})<f_{i^{\prime}}(\mathbf{x}^{(b)})$ [branke2008multiobjective].
@@
Solution $\mathbf{x}^{*}\in\mathcal{X}$ is considered Pareto optimal or non-dominated if there exists no other candidate solutions that can dominate $\mathbf{x}^{*}$. Consequently, Pareto Set (PS) and Pareto Front (PF) contain all the set of Pareto optimal solutions and their corresponding objective vectors [branke2008multiobjective].
@@ 2.2 Parametric Expensive Multi-Objective Optimization
In many real-world scenarios, EMOPs must be solved repeatedly under varying environmental [moo-gamut] or operational conditions [moo-qd]. For example, 3D model design considers optimizing geometric features to maintain the structural performance under varying external physical constraints [moo-loading]. These variations can be represented using continuous task parameters $\boldsymbol{\theta}\in\Theta\subseteq\mathbb{R}^{V}$, which defines a family of different EMOPs over the same decision space. This leads to the formulation of a parametric expensive multi-objective optimization problem (P-EMOP):
@@
where each task parameter $\boldsymbol{\theta}$ defines a distinct and expensive multi-objective problem instance $F(\mathbf{x},\boldsymbol{\theta})$. The ideal goal of P-EMOP is to approximate the Pareto Set across potentially infinite task parameters under limited evaluation budgets. We address this by learning an inverse model $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ that generates optimized solutions for any task parameter $\boldsymbol{\theta}$ and preference vector $\boldsymbol{\lambda}$ without expensive re-optimization. This represents a fundamental shift from per-task optimization to learning generalized solution mappings across the continuous parameter space. Here, $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot):\mathbb{R}^{M}\rightarrow\mathbb{R}$ is a scalarization function [^1] In this work, we instantiate $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot)$ for all methods using a recently proposed hypervolume scalarization function [hv-scalar]. characterized by preference vector $\boldsymbol{\lambda}$. We incorporate $\boldsymbol{\lambda}$ into the inverse model since the Pareto Set generally lies on a continuous manifold for non-trivial multi-objective optimization [moo-manifold], enabling $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ to output a single solution vector rather than a solution set.
Although (2) contains infinitely many tasks, these EMOPs share inter-task synergies due to their parametric relationship. Rather than treating tasks in isolation, we solve a finite sample $\{\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{K}\}$ as a multi-task optimization problem [wei], assuming tasks belong to a coherent family sharing solution components or objective landscape regularities [moo-benchmark]. This assumption, pioneered by multi-objective multifactorial optimization [moomfea] and commonly exploited in subsequent multi-task optimization works [mfea, moo-mfea2, moo-finv], enables more efficient solution search across the task parameter space through inter-task synergies, ultimately yielding higher-quality solutions for constructing the inverse model.
@@ 2.3 Gaussian Processes for Multi-Objective Optimization
In this paper, we adopt Gaussian Process (GP) [GPML] as surrogate models to efficiently solve expensive multi-objective optimization problems (EMOPs). A GP defines a distribution over functions, assuming that $f\sim\mathcal{GP}(\mu(\cdot),\kappa(\cdot,\cdot))$, where $\mu(\mathbf{x})=\mathbb{E}[f(\mathbf{x})]$ is the mean function and $\kappa(\mathbf{x},\mathbf{x}^{\prime})=Cov[f(\mathbf{x}),f(\mathbf{x}^{\prime})]$ is the kernel or covariance function. Given a dataset $\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{N}$, where each observation is corrupted by Gaussian noise $y^{(i)}=f(\mathbf{x}^{(i)})+\epsilon^{(i)},\epsilon^{(i)}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, the posterior at a test point $\mathbf{x}$ is a normal distribution $\mathcal{N}(\mu(\mathbf{x}),\sigma^{2}(\mathbf{x}))$ with:
@@
In EMOPs, Gaussian Processes (GPs) serve as probabilistic surrogates for approximating costly objective functions, enabling uncertainty-aware exploration using acquisition functions. Following the theoretically sound multi-objective Bayesian optimization (MOBO) framework in  [parego-2], in this paper, we adopt independent GP models for each objective and apply a scalarization-based acquisition function. In each iteration, we randomly sample a preference vector $\boldsymbol{\lambda}\in\Lambda$ that defines a trade-off among objectives via the scalarization function $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot)$, and optimize a scalarized acquisition function score, defined as:
@@ 3.3 Theoretical Analysis of Faster Convergence
For our theoretical analysis, we instantiate the acquisition functions $\boldsymbol{\alpha}(\mathbf{x})$ and $\boldsymbol{\widetilde{\alpha}}(\mathbf{x},\boldsymbol{\theta}_{k})$ using the Upper Confidence Bound (UCB) [andrea]. Specifically, for ST-MOBO, the acquisition function in (4) becomes:
@@
For an expensive multi-objective optimization problem with a total evaluation budget $T$, the goal is to obtain a set of solutions $\mathbf{X}_{T}\subset\mathcal{X}$ that approximates the true Pareto Set $\mathcal{X}^{*}$. Following the analysis technique [parego-2], this goal can be quantified using the notion of Bayes Regret, which evaluates the quality of the obtained solution set with respect to random preference vectors.
@@
Deriving an upper bound on Bayes regret is non-trivial [hv-scalar], since it measures the collective regret of the entire solution set rather than point-wise regret. Therefore, a surrogate of Bayes regret is introduced, termed as Cumulative Regret, which quantifies the total point-wise regret incurred across all evaluation rounds.
@@
In the theoretical study, we assume that each objective function $f_{m}(\cdot)$ follows a Gaussian distribution with marginal variances upper bounded by 11 and with independent noise variance upper bounded by $\sigma^{2}$ on a compact domain $\mathcal{X}\subseteq[0,1]^{D}$. Moreover, it is also assumed that scalarization functions $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot)$ is $J_{\boldsymbol{\lambda}}$-Lipschitz ($J_{\boldsymbol{\lambda}}\leq J,\forall\boldsymbol{\lambda}\in\Lambda$, $\Lambda$ being a bounded subset of a normed linear space) in the $\ell_{1}$-norm and monotonically increasing for each objective.
Theorem 1 (Regret Bounds for ST-MOBO, Theorem 1 in [parego-2]).
@@
where II stands for mutual information. Although Theorem 1 is stated for a single EMOP instance, it applies equally well to each of the $K$ tasks in isolation. For task $k$, let
@@
The detailed proof of Theorem 2 can be found in the Appendix.
@@ 3.4 Conditional Generative Model
Pure acquisition‐based search via PMT‐MOBO provides sample‐efficient guidance, but can steer sampling into low‐quality regions when $\mathcal{X}$ is large in (5) and budgets are tight [cbas, LFI]. To complement this, we learn an elite‐solution distribution via a conditional generative model. As illustrated in Figure 1, we let
@@
In this way, distribution‐aware sampling is seamlessly blended with uncertainty‐driven exploration. We instantiate this conditional generative model via either Variational Autoencoder (VAE) [vae] or Denoising Diffusion Probabilistic Model (DDPM) [ddpm] as follows.
@@
An encoder $q_{\phi_{e}}(z|\mathbf{x},\boldsymbol{c})$ approximates the posterior, and a decoder $p_{\phi_{d}}(\mathbf{x}|z,\boldsymbol{c})$ models the conditional likelihood. We learn $\phi=(\phi_{e},\phi_{d})$ by maximizing the variational lower bound [vae, cvae] on $\mathcal{D}$. During generative solution sampling, candidate solutions $\mathbf{x}^{\prime}$ can be obtained from the learned distribution as follows:
@@
with $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\hat{\beta}_{s})$. Thereafter, a neural net $\epsilon_{\phi}(\mathbf{x}_{t},t,\boldsymbol{c})$ is trained to predict the noise added at timestep $t$ given the noisy solution $\mathbf{x}_{t}$ and conditioning $\boldsymbol{c}$. We learn $\phi$ by minimizing the denoising objective following [ddpm]. The learned diffusion model parameterizes the conditional distribution of elite solutions as follows:

### v3

@@ Abstract
Many real-world applications require solving families of expensive multi-objective optimization problems (EMOPs) under varying operational conditions. This gives rise to parametric expensive multi-objective optimization problems (P-EMOPs) where each task parameter defines a distinct optimization instance. Current multi-objective Bayesian optimization methods have been widely used for finding finite sets of Pareto optimal solutions for individual tasks. However, P-EMOPs present a fundamental challenge: the continuous task parameter space can contain infinite distinct problems, each requiring separate expensive evaluations. This demands learning an inverse model that can directly predict optimized solutions for any task-preference query without expensive re-evaluation. This paper introduces a novel parametric multi-task multi-objective Bayesian optimizer that learns this inverse model by alternating between (1) acquisition-driven search leveraging inter-task synergies and (2) generative solution sampling via conditional generative models. This approach enables efficient optimization across related tasks and finally achieves direct solution prediction for unseen parameterized EMOPs without additional expensive evaluations. We theoretically justify the faster convergence by leveraging inter-task synergies through task-aware Gaussian processes. Meanwhile, based on that, empirical studies of our optimizer and inverse model in synthetic and real-world benchmarks further verify the effectiveness of the proposed generative alternating framework.
@@ 1 Introduction
Expensive multi-objective optimization problems (EMOPs) are crucial in black-box optimization, aiming to identify optimal trade-off solutions among multiple conflicting criteria under limited evaluation budgets [27, 38]. Due to their focus on conflicting objective functions and costly evaluations, EMOP formulations have been applied to a plethora of real-world applications encompassing materials discovery [36, 29], drug discovery [47, 66], and robotics [54, 24].
@@
To efficiently identify trade-off solutions, surrogate modeling has become essential for reducing the cost of evaluating expensive objectives. Among these, multi-objective Bayesian optimization (MOBO) stands out as the leading strategy. MOBO methods have been widely used for identifying finite sets of Pareto optimal solutions by scalarizing multiple objectives into single objectives [27, 44, 41] or by optimizing metrics such as expected hypervolume improvement [14, 11, 12] and uncertainty reduction [42, 20, 53, 3]. To move beyond these finite solution‑set approximations and directly model the entire trade‑off surface, recent efforts have extended MOBO to continuous Pareto‑manifold learning [32, 34, 31].
@@
However, in many real-world scenarios, optimization problems arise not as isolated instances but as families of related tasks under varying operational conditions. For example, turbine components must be optimized across different wind speeds [35], and robotic control policies must be optimized for diverse terrains [55] or damage conditions [10]. These cases create fundamental challenges: the continuous range of operational conditions forms infinitely many related optimization problems, each demanding separate, costly evaluations under current MOBO methods. As operational conditions frequently shift [39], repeatedly performing full optimizations to evaluate trade-offs across multiple scenarios becomes prohibitively expensive and computationally infeasible. Therefore, handling families of related EMOPs requires new approaches that avoid excessive evaluation costs. To our best knowledge, little existing work has systematically addressed this parametric setting.
@@
This paper proposes a novel parametric multi-objective optimizer to address this challenge by learning an inverse model. This inverse model can directly predict optimized solutions for any given task parameter, thereby bypassing the repeated costly evaluations mentioned above. Our approach alternates between two complementary phases: (1) acquisition-driven search guided by task-aware Gaussian processes, which explicitly leverages inter-task correlations to efficiently explore the solution space, and (2) generative solution sampling via conditional generative models, which learns high-quality solution distributions conditioned on task parameters and preferences. This dual strategy not only significantly enhances optimization efficiency across multiple related tasks but also finally builds an inverse model for predicting optimized solutions to previously unseen EMOPs without additional expensive evaluations.
@@
  • We propose a novel inverse generative model to predict optimized solutions to each task query within the continuous task parameter space, thereby addressing the challenge of solving families of EMOPs.
@@ 2.1 Expensive Multi-Objective Optimization
Solution $\mathbf{x}^{(a)}\in\mathcal{X}$ is considered Pareto dominate another solution $\mathbf{x}^{(b)}\in\mathcal{X}$ if $\forall i\in[M]$, $f_{i}(\mathbf{x}^{(a)})\leq f_{i}(\mathbf{x}^{(b)})$ and $\exists\,i^{\prime}\in[M]$ such that $f_{i^{\prime}}(\mathbf{x}^{(a)})<f_{i^{\prime}}(\mathbf{x}^{(b)})$ [4].
@@
Solution $\mathbf{x}^{*}\in\mathcal{X}$ is considered Pareto optimal or non-dominated if there exists no other candidate solutions that can dominate $\mathbf{x}^{*}$. Consequently, Pareto Set (PS) and Pareto Front (PF) contain all the set of Pareto optimal solutions and their corresponding objective vectors [4].
@@ 2.2 Parametric Expensive Multi-Objective Optimization
In many real-world scenarios, EMOPs must be solved repeatedly under varying environmental [35] or operational conditions [43]. For example, 3D model design considers optimizing geometric features to maintain the structural performance under varying external physical constraints [45]. These variations can be represented using continuous task parameters $\boldsymbol{\theta}\in\Theta\subseteq\mathbb{R}^{V}$, which defines a family of different EMOPs over the same decision space. This leads to the formulation of a parametric expensive multi-objective optimization problem (P-EMOP):
@@
where each task parameter $\boldsymbol{\theta}$ defines a distinct and expensive multi-objective problem instance $F(\mathbf{x},\boldsymbol{\theta})$. The ideal goal of P-EMOP is to approximate the Pareto Set across potentially infinite task parameters under limited evaluation budgets. We address this by learning an inverse model $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ that generates optimized solutions for any task parameter $\boldsymbol{\theta}$ and preference vector $\boldsymbol{\lambda}$ without expensive re-optimization. This represents a fundamental shift from per-task optimization to learning generalized solution mappings across the continuous parameter space. Here, $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot):\mathbb{R}^{M}\rightarrow\mathbb{R}$ is a scalarization function [^1] In this work, we instantiate $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot)$ for all methods using a recently proposed hypervolume scalarization function [65]. characterized by preference vector $\boldsymbol{\lambda}$. We incorporate $\boldsymbol{\lambda}$ into the inverse model since the Pareto Set generally lies on a continuous manifold for non-trivial multi-objective optimization [37], enabling $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ to output a single solution vector rather than a solution set.
Although (2) contains infinitely many tasks, these EMOPs share inter-task synergies due to their parametric relationship. Rather than treating tasks in isolation, we solve a finite sample $\{\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{K}\}$ as a multi-task optimization problem [58], assuming tasks belong to a coherent family sharing solution components or objective landscape regularities [61]. This assumption, pioneered by multi-objective multifactorial optimization [18] and commonly exploited in subsequent multi-task optimization works [19, 2, 57] and recent parametric multi-objective optimization studies [7, 8], enables more efficient solution search across the task parameter space through inter-task synergies, ultimately yielding higher-quality solutions for constructing the inverse model.
@@ 2.3 Gaussian Processes for Multi-Objective Optimization
In this paper, we adopt Gaussian Process (GP) [59] as surrogate models to efficiently solve expensive multi-objective optimization problems (EMOPs). A GP defines a distribution over functions, assuming that $f\sim\mathcal{GP}(\mu(\cdot),\kappa(\cdot,\cdot))$, where $\mu(\mathbf{x})=\mathbb{E}[f(\mathbf{x})]$ is the mean function and $\kappa(\mathbf{x},\mathbf{x}^{\prime})=Cov[f(\mathbf{x}),f(\mathbf{x}^{\prime})]$ is the kernel or covariance function. Given a dataset $\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{N}$, where each observation is corrupted by Gaussian noise $y^{(i)}=f(\mathbf{x}^{(i)})+\epsilon^{(i)},\epsilon^{(i)}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, the posterior at a test point $\mathbf{x}$ is a normal distribution $\mathcal{N}(\mu(\mathbf{x}),\sigma^{2}(\mathbf{x}))$ with:
@@
In EMOPs, Gaussian Processes (GPs) serve as probabilistic surrogates for approximating costly objective functions, enabling uncertainty-aware exploration using acquisition functions. Following the theoretically sound multi-objective Bayesian optimization (MOBO) framework in  [41], in this paper, we adopt independent GP models for each objective and apply a scalarization-based acquisition function. In each iteration, we randomly sample a preference vector $\boldsymbol{\lambda}\in\Lambda$ that defines a trade-off among objectives via the scalarization function $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot)$, and optimize a scalarized acquisition function score, defined as:
@@ 3.3 Theoretical Analysis of Faster Convergence
For our theoretical analysis, we instantiate the acquisition functions $\boldsymbol{\alpha}(\mathbf{x})$ and $\boldsymbol{\widetilde{\alpha}}(\mathbf{x},\boldsymbol{\theta}_{k})$ using the Upper Confidence Bound (UCB) [49]. Specifically, for ST-MOBO, the acquisition function in (4) becomes:
@@
For an expensive multi-objective optimization problem with a total evaluation budget $T$, the goal is to obtain a set of solutions $\mathbf{X}_{T}\subset\mathcal{X}$ that approximates the true Pareto Set $\mathcal{X}^{*}$. Following the analysis technique [41], this goal can be quantified using the notion of Bayes Regret, which evaluates the quality of the obtained solution set with respect to random preference vectors.
@@
Deriving an upper bound on Bayes regret is non-trivial [65], since it measures the collective regret of the entire solution set rather than point-wise regret. Therefore, a surrogate of Bayes regret is introduced, termed as Cumulative Regret, which quantifies the total point-wise regret incurred across all evaluation rounds.
@@
In the theoretical study, we assume that each objective function $f_{m}(\cdot)$ follows a Gaussian distribution with marginal variances upper bounded by $1$ and with independent noise variance upper bounded by $\sigma^{2}$ on a compact domain $\mathcal{X}\subseteq[0,1]^{D}$. Moreover, it is also assumed that scalarization functions $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot)$ is $J_{\boldsymbol{\lambda}}$-Lipschitz ($J_{\boldsymbol{\lambda}}\leq J,\forall\boldsymbol{\lambda}\in\Lambda$, $\Lambda$ being a bounded subset of a normed linear space) in the $\ell_{1}$-norm and monotonically increasing for each objective.
Theorem 1 (Regret Bounds for ST-MOBO, Theorem 1 in [41]).
@@
where $I$ stands for mutual information. Although Theorem 1 is stated for a single EMOP instance, it applies equally well to each of the $K$ tasks in isolation. For task $k$, let
@@
The detailed proof of Theorem 2 can be found in the Appendix A.
@@ 3.4 Conditional Generative Model
Pure acquisition‐based search via PMT‐MOBO provides sample‐efficient guidance, but can steer sampling into low‐quality regions when $\mathcal{X}$ is large in (5) and budgets are tight [5, 62]. To complement this, we learn an elite‐solution distribution via a conditional generative model. As illustrated in Figure 1, we let
@@
In this way, distribution‐aware sampling is seamlessly blended with uncertainty‐driven exploration. We instantiate this conditional generative model via either Variational Autoencoder (VAE) [25] or Denoising Diffusion Probabilistic Model (DDPM) [21] as follows.
@@
An encoder $q_{\phi_{e}}(z|\mathbf{x},\boldsymbol{c})$ approximates the posterior, and a decoder $p_{\phi_{d}}(\mathbf{x}|z,\boldsymbol{c})$ models the conditional likelihood. We learn $\phi=(\phi_{e},\phi_{d})$ by maximizing the variational lower bound [25, 48] on $\mathcal{D}$. During generative solution sampling, candidate solutions $\mathbf{x}^{\prime}$ can be obtained from the learned distribution as follows:
@@
with $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\hat{\beta}_{s})$. Thereafter, a neural net $\epsilon_{\phi}(\mathbf{x}_{t},t,\boldsymbol{c})$ is trained to predict the noise added at timestep $t$ given the noisy solution $\mathbf{x}_{t}$ and conditioning $\boldsymbol{c}$. We learn $\phi$ by minimizing the denoising objective following [21]. The learned diffusion model parameterizes the conditional distribution of elite solutions as follows: